# Eco-evolutionary dynamics of cooperative antimicrobial resistance in a population of fluctuating volume and size

### v1

@@ 4.3 Theory for the fixation-coexistence diagrams in fluctuating environments
We notice however that in the presence of EV each panel of figure 4 is coded in two colours, indicating a phase of fast fixation/extinction and another of long-lived coexistence. In fact, due to EV, long-lived coexistence is possible even when fast fixation of $S$ and $R$ is most likely. In figure 4(a),(d) this leads to the red and blue dominated diagrams to also contain a dark/black area for $\delta$ close to $1$ (bias of $K$ towards $K_{+}$). This can be qualitatively explained by noting that, as $K$ is biased towards $K_{+}$, the population size is essentially constant and large, $N\approx K_{+}$, and does not admit fast fixation of either strain. Similarly, EV can allow for fast fixation even when long-lived coexistence is expected, as found in figure 4(b), where the black-dominated diagram (due to $x_{th}\approx x^{*}_{th}$) contains a coloured “cloud”. Here, fast fixation events occur for low values of $\nu$ and $\delta\leq 0$, when $N\approx K_{-}$ for long periods, and the small population size ($K_{-}=100$) facilitates early extinction of either $R$ or $S$, see figure 5.
@@
In the aim of understanding the evolution of AMR, we are particularly interested in finding the conditions favouring the early extinction of $R$, and hence the fast fixation of $S$, since this corresponds to the most favourable environmental conditions for the early eradication of AMR [HNARM23a]. Here, the best conditions for AMR eradication appear as red areas in the diagrams of figure 4: these are low values of $x_{\text{th}}$ and $\delta$ not too close to $1$, see figure 4(a). Moreover, as a signature of the influence of EV, discussed above, we find that AMR eradication is also possible for intermediate values of $x_{\text{th}}\gtrsim x_{\text{th}}^{*}$, see reddish area in figure 4(b) and inset of figure 5(b).
@@ 5 Conclusions
Microorganisms live in endlessly changing environments, and experience conditions that often fluctuate between mild and harsh. This environmental variability (EV) often affects the amount of nutrients and toxins in a community, and thus shapes the eco-evolutionary dynamics of microbial communities. EV often crucially influences the ability of species to coexist and cooperate, a feature that is key to better understand the evolution of antimicrobial resistance (AMR), whose rise is a global societal threat [O’N16, Col22]. In this context, a central question concerns how EV affects the coexistence of microbes resistant and sensitive to antimicrobial drugs. It is well established that certain resistant cells are able to inactivate antimicrobials and, under certain conditions, share the protection against the drugs with the entire microbial community. This form of AMR can be regarded as cooperative behaviour. In this work, we have studied an idealised model of cooperative AMR where sensitive and resistant cells compete for the same resources in a microbial community of fluctuating size, in the presence of a biostatic drug, and subject to demographic fluctuations and environmental variability (EV). The latter is modelled by a binary carrying capacity randomly switching between high and low values (mild and harsh conditions, respectively), and is responsible for the time-variation of the community volume and size: the total amount of nutrients and toxin in the community fluctuates, but their concentration is kept fixed (time-changing volume). When the concentration of resistant microbes exceeds a fixed concentration cooperation threshold, their protection, provided by drug-inactivating enzymes, is shared with sensitive cells at no metabolic cost. However, when the resistant cells concentration is below the cooperation threshold, only resistant microbes benefit from the protection offered by the enzymes. In this setting, the evolutionary dynamics of AMR can be viewed as the eco-evolutionary dynamics of a public good game in a fluctuating environment for a microbial community of time-varying volume. In a large population under static environment, fluctuations are negligible and both strains always coexist. Moreover, in the absence of EV, the fate of a community of finite size is chiefly determined by the value of the concentration cooperation threshold $x_{\text{th}}$ relative to a critical value $x^{*}_{\text{th}}$ that we have analytically determined. In particular, when the concentration cooperation threshold is close to 0, which means that resistant microbes are very efficient at generating drug-inactivating enzymes, resistant cells are likely to become extinct. On the other hand, when resistant microbes are inefficient at inactivating the drug, the cooperation threshold is close to 1, and the resistant strain takes over the entire microbial population. Otherwise, when the cooperation threshold is close to the critical value $x^{*}_{\text{th}}$, a long-lived coexistence of the strains is expected to set in.
@@
It is worth noticing that in [HNARM23a] we considered a related model with environmental fluctuations arising at constant volume, i.e. a form of EV with time-varying concentration of nutrients and toxins, which is characterised by fixation-coexistence diagrams of three phases (fixation of each strain and long-lived coexistence), and a fluctuation-driven AMR eradication mechanism. Hence, the findings reported in this work complement those of [HNARM23a] and give us a broader perspective on the joint influence of environmental and demographic fluctuations on the evolution of cooperative AMR, paving the way for numerous possible applications. We also believe that the analytical methods devised in this work, and their generalisations, shall be applicable to describe the eco-evolutionary dynamics of a broad class of systems.
@@
Contributor roles taxonomy by CRediT [${}^{+}$15].
@@ Appendix A Simulation methods
We have studied the full eco-evolutionary dynamics of the model by carrying out extensive exact stochastic simulations following the Next Reaction Method of references [GB00, And07], and proceeding as in [AHNRM23, HNARM23a], where details and code are presented in the supplemental material [HNARM23b]. The simulation data and codes that we have generated are electronically available, alongside comments, in [HNAM23].
@@
Appendix B Mean Coexistence Time (MCT)
In this section we derive the approximate, analytical expression for the mean fixation or extinction time at fixed total population $N$, i.e. the mean coexistence time (MCT) $\langle\tau_{N}\rangle$. Supported by the assumption $\langle\tau\left(x_{0}|x_{\text{th}},s,a,N\right)\rangle\simeq\langle\tau\left(x_{0}=x_{\text{th}}|x_{\text{th}},s,a,N\right)\rangle$ of section 3.2, we focus on the case $N_{R}^{0}=N_{\text{th}}\equiv x_{\text{th}}N$ to simplify the main equation (12). Note that, in this case, the last double summation has powers of $\gamma=(1-a)/(1-s)<1$, while the first double summation also has additional summands with high powers of $\gamma=1/(1-s)>1$. Hence, the last double summation is negligible with respect to the first one. Furthermore, if we split the inner summation of the first double summation at $n=N_{\text{th}}$ we obtain
@@ Appendix B Mean Coexistence Time (MCT)
Appendix C Numerical marginalisation of an arbitrary function over the N-PDMP
@@
The marginal stationary probability density function of $N$-PDMP is [Hor84, HL06, Ben06, RDL11, WFM17, WFM18, TWAM20, AHNRM23]

### v2

@@ 4.3 Theory for the fixation-coexistence diagrams in fluctuating environments
We notice however that in the presence of EV each panel of figure 4 is coded in two colours, indicating a phase of fast fixation/extinction and another of long-lived coexistence. In fact, due to EV, long-lived coexistence is possible even when fast fixation of $S$ and $R$ is most likely. In figure 4(a),(d) this leads to the red and blue dominated diagrams to also contain a dark/black area for $\delta$ close to $1$ (bias of $K$ towards $K_{+}$). This can be qualitatively explained by noting that, as $K$ is biased towards $K_{+}$, the population size is essentially constant and large, $N\approx K_{+}$, implying that demographic fluctuations are weak, which prevents fast fixation of either strain. Similarly, EV can allow for fast fixation even when long-lived coexistence is expected, as found in figure 4(b), where the black-dominated diagram (due to $x_{th}\approx x^{*}_{th}$) contains a coloured “cloud”. Here, fast fixation events occur for low values of $\nu$ and $\delta\leq 0$, when $N\approx K_{-}$ for long periods, and the small population size ($K_{-}=100$) enforces large demographic fluctuations that facilitate early extinction of either $R$ or $S$, see figure 5.
@@
In the aim of understanding the evolution of AMR, we are particularly interested in finding the conditions favouring the early extinction of $R$, and hence the fast fixation of $S$, since this corresponds to the most favourable environmental conditions for the early eradication of AMR [32]. Here, the best conditions for AMR eradication appear as red areas in the diagrams of figure 4: these are low values of $x_{\text{th}}$ and $\delta$ not too close to $1$, see figure 4(a). Moreover, as a signature of the influence of EV, discussed above, we find that AMR eradication is also possible for intermediate values of $x_{\text{th}}\gtrsim x_{\text{th}}^{*}$, see reddish area in figure 4(b) and inset of figure 5(b).
@@ 5 Conclusions
Microorganisms live in endlessly changing environments, and experience conditions that often fluctuate between mild and harsh. This environmental variability (EV) often affects the amount of nutrients and toxins in a community, and thus shapes the eco-evolutionary dynamics of microbial communities. EV often crucially influences the ability of species to coexist and cooperate, a feature that is key to better understand the evolution of antimicrobial resistance (AMR), whose rise is a global societal threat [50, 51]. In this context, a central question concerns how EV affects the coexistence of microbes resistant and sensitive to antimicrobial drugs. It is well established that certain resistant cells are able to inactivate antimicrobials and, under certain conditions, share the protection against the drugs with the entire microbial community. This form of AMR can be regarded as cooperative behaviour. In this work, we have studied an idealised model of cooperative AMR where sensitive and resistant cells compete for the same resources in a microbial community of fluctuating size, in the presence of a biostatic drug, and subject to demographic fluctuations and environmental variability (EV). The latter is modelled by a binary carrying capacity randomly switching between high and low values (mild and harsh conditions, respectively), and is responsible for the time-variation of the community volume and size: the total amount of nutrients and toxin in the community fluctuates, but their concentration is kept fixed (time-changing volume). When the concentration of resistant microbes exceeds a fixed concentration cooperation threshold, their protection, provided by drug-inactivating enzymes, is shared with sensitive cells at no metabolic cost. However, when the resistant cells concentration is below the cooperation threshold, only resistant microbes benefit from the protection offered by the enzymes. In this setting, the evolutionary dynamics of AMR can be viewed as the eco-evolutionary dynamics of a public good game in a fluctuating environment for a microbial community of time-varying volume. In a large population under static environment, fluctuations are negligible and both strains always coexist. Moreover, in the absence of EV, the fate of a community of finite size is chiefly determined by the value of the concentration cooperation threshold $x_{\text{th}}$ relative to a critical value $x^{*}_{\text{th}}$ that we have analytically determined. In particular, when the concentration cooperation threshold is close to 0, which means that resistant microbes are very efficient at generating drug-inactivating enzymes, resistant cells are likely to become extinct. On the other hand, when resistant microbes are inefficient at inactivating the drug, the cooperation threshold is close to 1, and the resistant strain takes over the entire microbial population. Otherwise, when the cooperation threshold is close to the critical value $x^{*}_{\text{th}}$, a long-lived coexistence of the strains is expected to set in.
@@
It is worth noticing that in [32] we considered a related model with environmental fluctuations arising at constant volume, i.e. a form of EV with time-varying concentration of nutrients and toxins, which is characterised by fixation-coexistence diagrams of three phases (fixation of each strain and long-lived coexistence), and a fluctuation-driven AMR eradication mechanism. Hence, the findings reported in this work complement those of [32] and give us a broader perspective on the joint influence of environmental and demographic fluctuations on the evolution of cooperative AMR, paving the way for numerous possible applications. We also believe that the analytical methods devised in this work, and their generalisations, shall be applicable to describe the eco-evolutionary dynamics of a broad class of systems.
@@
Contributor roles taxonomy by CRediT [98].
@@ Appendix A Simulation methods
We have studied the full eco-evolutionary dynamics of the model by carrying out extensive exact stochastic simulations following the Next Reaction Method of references [85, 86], and proceeding as in [13, 32], where details and code are presented in the supplemental material [99]. The simulation data and codes that we have generated are electronically available, alongside comments, in [100].
@@
Appendix B Fixation probability
For the Moran model and a population of constant size $N$, the exact general expression of the $R$ fixation probability depending on the initial number of $R$ cells is [72, 73, 33, 76, 74]
$\phi_{N}\left(N_{R}^{0}\right)=\frac{1+\sum_{k=1}^{N_{R}^{0}-1}\prod_{i=1}^{k}\gamma\left(i,N\right)}{1+\sum_{k=1}^{N-1}\prod_{i=1}^{k}\gamma\left(i,N\right)},\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \gamma\left(N_{R},N\right)\equiv\frac{\widetilde{T}^{-}_{R}\left(N_{R},N\right)}{\widetilde{T}^{+}_{R}\left(N_{R},N\right)}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ N_{R}^{0}=1,2,...,N,$ (S1)
where $\widetilde{T}^{\pm}_{R}\left(N_{R},N\right)$ are defined in the main equation (9) and the function $\gamma\left(N_{R},N\right)$ fully determines the fixation probability. When $N$ is sufficiently large, the initial fraction of resistant cells $x_{0}\equiv N_{R}^{0}/N$ can be approximated to a continuous variable. The fixation probability can thus be regarded as a continuous function of $x_{0}$, and it is convenient to write $\phi_{N}\left(N_{R}^{0}=Nx_{0}\right)$ as $\phi_{N}\left(x_{0}\right)$.
As discussed in main section 3.2, when $x_{0}$ is not too close to 00 or $1$, we can assume that the fixation/extinction of a strain occurs from $x_{0}=x_{\text{th}}$, i.e., $\phi_{N}(x_{0})\simeq\phi_{N}\left(x_{0}=x_{\text{th}}\right)$. In this case, the fixation probability can be simplified to the following exact expression [32]:
$\phi_{N}(x_{\text{th}})=\frac{1-\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}}{1-\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}+\frac{a-s}{s(1-a)}\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}\left[\left(\frac{1}{1-s}\right)^{(1-x_{\text{th}})N}-1\right]}.$ (S2)
The compact approximation of main equation (10) has been obtained by neglecting $(1-a)^{x_{\text{th}}N}/(1-s)^{x_{\text{th}}N}\ll 1$ and $(1-s)^{(1-x_{\text{th}})N}\ll 1$.
Appendix C Mean Coexistence Time (MCT)
The exact expression of the MCT for the Moran model in a fixed population of fixed size $N$ is well-known [72, 73, 33, 76, 74], and can here be written as
$\langle\tau_{N}\left(N^{0}_{R}\right)\rangle=\phi_{N}(N^{0}_{R})\sum_{k=N_{R}^{0}}^{N-1}\sum_{n=1}^{k}\frac{\prod_{m=n+1}^{k}\gamma\left(m,N\right)}{\widetilde{T}^{+}_{R}\left(n,N\right)}-\left[1-\phi_{N}(N^{0}_{R})\right]\sum_{k=1}^{N_{R}^{0}-1}\sum_{n=1}^{k}\frac{\prod_{m=n+1}^{k}\gamma\left(m,N\right)}{\widetilde{T}^{+}_{R}\left(n,N\right)},$ (S3)
where the subscript indicates that the population size $N$ is fixed (here $N=K_{0}$). The function $\phi_{N}(N^{0}_{R})$ is the $R$ fixation probability given by equation (S1), where the functions $\widetilde{T}^{+}_{R}\left(N_{R},N\right)$ and $\gamma\left(N_{R},N\right)$ are also defined.
Supported by the assumption $x_{0}\simeq x_{\text{th}}$ of section 3.2, we focus on the case $N_{R}^{0}=N_{\text{th}}\equiv x_{\text{th}}N$ to simplify equation (S3). Note that, in this case, the last double summation has powers of $\gamma=(1-a)/(1-s)<1$, while the first double summation also has additional summands with high powers of $\gamma=1/(1-s)>1$. Hence, the last double summation is negligible with respect to the first one. Furthermore, if we split the inner summation of the first double summation at $n=N_{\text{th}}$ we obtain
@@ Appendix C Mean Coexistence Time (MCT)
Appendix D Numerical marginalisation of an arbitrary function over the N-PDMP
@@
The marginal stationary probability density function of $N$-PDMP is [95, 80, 79, 81, 11, 12, 40, 13]